OpenStudy (anonymous):

calculate the standard deviation for the following data set: 16,35,25,29,26,30,42,18

OpenStudy (amistre64):

find the mean (xbar) subtract the mean from every score; x - xbar = a square the differences; a^2 divide a^2 by (n-1) where n is the number of data points = b sqrt(b) is teh standard deviation

OpenStudy (amistre64):

16,18,25,26,(mean),29,30,35,42 mean = 27.5

OpenStudy (anonymous):

A.7.9 B.8.5 C.26 D.28

OpenStudy (amistre64):

16-27.5 = -11.5 18-27.5 = - 9.5 25-27.5 = - 2.5 26-27.5 = - 1.5 29-27.5 = 1.5 30-27.5 = 2.5 35-27.5 = 7.5 42-27.5 = 14.5 ----------------- 1

OpenStudy (amistre64):

forget the 1 lol

OpenStudy (amistre64):

square then add...

OpenStudy (anonymous):

:/

OpenStudy (amistre64):

506 is the total of the squres

OpenStudy (amistre64):

sqrt(506/7) = standard deviation... ....8.50

OpenStudy (amistre64):

id go with B